OpenStudy (anonymous):

A positive integer is divided by a smaller positive integer. The integer part of the quotient is 9 and the remainder is 1. The sum of the two integers is 71. Find the two integers. Write the two answers separated by a comma with no spaces.

OpenStudy (rogue):

x + y = 71 x/y = 9 R 1 (x-1)/y = 9 x - 1 = 9y x = 9y + 1 x + y = 71 (9y + 1) + y = 71 10 y + 1 = 71 10y = 70 y = 7 x + y = 71 x + 7 = 71 x = 64 x = 64, y = 7